A Sample Foils Document

Jim Hafner IBM Research Division Almaden Research Center

        hafner@almaden.ibm.com

September 15, 1999

Abstract

This is where an abstract might go if you want one. There is usually not a lot of room for much here.

– Typeset by FoilTEX – Is this for a restricted audience?

Jim Hafner 1998/08/15

Colors

FoilTEX (version 2) is fully integrated with L^{A}TEX 2ε so that the supported color package (part of the graphics package) is the preferred way to use colors.

For examples, see the graphics package documentation.

You can still use the old colordvi package that comes with Rokicki’s dvips, but some things won’t work exactly as expected (except in compatibility mode).

Jim Hafner 1998/08/15

Itemize

• This is the first level of an itemize.

– Here we jump to second level

∗ Even third level (and there is a fourth level as well).

∗ The second item at third level.

– The second item at second level.

– A third item at second level.

• The second item at first level.

• A third item at first level.

Note that we have turned off the logo on this page. It returns on the next page.

Jim Hafner 1998/08/15

This should be a new foil with no header, followed by a quote:

… it’s a good idea to make your input file as easy to read as possible.

and some enumerating:

1. this is enumerated

2. this is also enumerated

Above, we used \emph{new foil} instead of

   the old {\em new foil\/}!

On the following page we decided to stop the headers from appearing and move the date to the footline.

Fonts

• This shows italics, slanted, boldface, typewriter, roman, and small caps.

• Unslanted emphasize and slanted emphasize.

• \textrm means roman and \textsf means sans serif.

• size changing from tiny, scriptsize, footnotesize, small, normalsize, large,Large,LARGE,huge,

and Huge.

• 12pt is the smallest preloaded.

• 43pt is the largest preloaded in compatibility mode.

• 51pt is the largest preloaded in normal mode.

– Typeset by FoilTEX – Is this for a restricted audience? 1998/08/15 4

Special Characters and Accents

• Here is a list of accents:

– ò, ó, ô, ö, õ, ō, ȯ, ŏ, ǒ, ő, oo, o̧, o., ō.

• First in paragraph mode (with \copyright):

– †, ‡, §, ¶, £, 0123456789, \begin{matrix}c \\ ◯ \end{matrix}

• Then in math mode: (numerals are different!) – †, ‡, §, ¶, £, 0123456789

• Here are more non-english language symbols: – œ, Œ, æ, Æ, å, Å, ø, Ø, ł, Ł, ß, ¿, ¡

• TEX’s special symbols: #, $, %, &, , {, }.

Some Mathematics

F ⋯ \frac{x + y}{1 + \frac{y}{z+1}} = \sqrt{x + y} × ^{n} \sqrt{2}

Here are some funny math symbols (we needed the latexsym package for a couple of these):

∀∃♭♮♯∂^{}ℜ̸℧ȷℵ ∞◇♣♢♠♡ℓ

⋂ ⋃ ⋁ ⋀ ⨀ ⨂ ⨁ ⨄ ∑ ∏ ∐ ∫ ∮

∑_{i=1}^{n} x_{i} = ∫_{0}^{1} f[x]dx.

and gcd(m, n) and x ≡ y (mod a + b).

– Typeset by FoilTEX – Is this for a restricted audience? 1998/08/15 6

More Math: arrays

\begin{matrix}a + b + c & uv & x − y & 27 \\ a + b & u + v & z & 134 \\ a & 3u + vw & xyz & 2, 978 \end{matrix}

( \begin{matrix}| & \begin{matrix}x_{11} & x_{12} \\ x_{21} & x_{22} \end{matrix} & | \\ & y \\ & z \end{matrix} )

x = \{ \begin{matrix}y & if y > 0 \\ z + y & otherwise \end{matrix}

– Typeset by FoilTEX – Is this for a restricted audience? 1998/08/15 7

 More Math: equation and array

Numbering

Here is a numbered equation

E = mc^{2}  (1)

and a numbered array

x = 17y  (2)
y > a + b + c + d + e + f +
k + l + m + n + o + p  (3)

More math accents: â, ǎ, ă, á, à, ã, ā, a⃗, ȧ, ä. Over and underline:

\overline{\overline{x}^{2} + 1}  \overbrace{a + \underbrace{b + c}_{25} +d}^{16α}

Bold Mathematics

The boldequation environment, with numbering yields

2\sqrt{x}Π^{y} ∼ π × x  (4)

and without numbering yields

2\sqrt{x}Π^{y} ∼ π × x

We can reference bold equations like (4).

There is also \mathbf and \bm in the middle of a formula, with

\mathbf{a +} a + \bm{a+x\pi-\rho} -\rho

a+a + a + xπ − ρ − ρ

Note the difference between the two bold “a” in result.

Theorem and Proof Environments

Theorem 1. [TUG’92] There are some TEX tools that are easier to use than others. This theorem is numbered and has an optional acknowledgement.

Corollary. This obvious corollary is not numbered because it uses the *-form.

Proof. The details of the proof are left to the reader. Note that the environment names are case sensitive. □

We put a header back on the next foil to see that it is correctly rotated.

I’m rotated!

Rotated Foils

This entire page will (should?) be rotated if we declared the \documentclass option dvips. In the other cases, this is not supported and the user is warned.

The next foil will return to normal. Only foils that begin with \rotatefoilhead will be rotated. If L^{A}TEX needs to split a rotated foil into two foils, both will be rotated.

Tables and Figures

Here is a short table:

First stuff Second stuff Third Stuff
foo  bar  bug
foofoo  barbar  bugbug

Table 1: This is the first table.

The above table is Table 1. It is a “nonfloat”, since it doesn’t float at all, but appears right where it was placed in the document.

picture Environments

[figure: gnat]

Figure 1: Isn’t this a pretty picture?

\thicklines would be nice here but the picture doesn’t look as good that way (why?). We should use the graphics package to load our graphics pictures here. Note that this Figure doesn’t float!

Marginal Notes, Footnotes and Citations

A marginal note^{1} is made with the \marginpar command, having the text as its argument. The note

This is a mar- ginal note

is placed in the margin, its first line even with the line of the text containing the command^{2}.

This is shows frameboxes at work. We can even cite references like [1] and [2].

References

[1] Rocky and Bullwinkle, Open problems, in Mr. Know-it-all’s Rock Encyclopedia.

[2] Bullwinkle, Getting things out of hats, Annals 1 (1990BC), 1–2.

1This is a footnote.

2This is a second footnote.

– Typeset by FoilTEX – Is this for a restricted audience? 1998/08/15 14

Other Features

• \raggedright can be used in the preamble to get this effect throughout as we did on this page.

• Access to the AMS Fonts symbols and Euler fonts can easily be obtained with the amslatex (or is it amsmath) packages. You might not get fonts at the largest sizes however.

• PostScript fonts can used just by adding any of the package files from the PSNFSS package.